Can you please help me write the folling equation in slope-intercept form 2x+5=3y.
Answer by tutorcecilia(2144)

2x+5=3y [re-write in slope-intercept form by solving for y]
y=mx+b [slope-intercept formulat]
2x/3+5/3=3y/3
2/3x+5/3=y or y=2/3x+5/3 [slope-intercept form]